OpenStudy (amenah8):

Find the local maximum and local minimum for each if they exist and identify the intervals when the function is increasing AND decreasing f(x)=(1/3)x^3 - 3x^2 + 8x

OpenStudy (irishboy123):

@Amenah8 what is y' ?????

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{ d }{dx} x^n=nx^{n-1} }\) that is the "Power Rule". Apply this rule to differentiate each term individually (and thus you will differentiate the function).